Operations Report – February 21st

Crew 189 Operations Report 21/02/2018
SOL: 3

Name of person filing report: Gabriel PAYEN

Non-nominal systems: Pipe leak, Fans on backpack 2 & 6, roof heater, light bulb in the bathroom, hesitation about burnable/unburnable trashes.

-Pipe :
The pipe in the low desk still leaks

-Roof Heater :
It’s still difficult to control the heater

-Backpacks :
All engineering checks were nominals, but during the EVA fans of backpack 2 and 6 turned out of service and both helmets were very foggy.
Backpacks 2 and 6 joined backpacks 3 and 4 to the list of backpacks the crew need to fix. The crew was supposed to fix them today but the we were all very busy. I hope we’ll begin tomorrow.

-Light bulb :
At 5:00 PM, a crew member was unable to light on the bulb in the bathroom. It seems that it is broken.

-Trashes :
During the EVA engineering check, the EVA members had to but five wastes bags into the outside trashes but “Burnable” and “Unburnable” weren’t indicated. Therefore, they randomly put the wastes into the trashes, which is quite pity considering we do a selective sorting for wastes.

Generator (hours run): Turned of at 12:20 and turned on at 18:08

Solar— SOC % (Before generator is run at night) –

Diesel – 70 %

Propane – 39 psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 7 gallons.

Water (trailer) – 0 gallons.

Water (static) – 300 gallons

Trailer to Static Pump used – NO

Water (loft) – Static to Loft Pump used – YES

Water Meter: 131 293.2

Toilet tank emptied: YES

ATV’s Used: 350.1 350.2

Oil Added : No

ATV Fuel Used: Gals

Hours the ATVs were Used today: 3 hours

Notes on ATVs: Because of the cold night, it took a few minutes to turn them on.

Deimos rover used: NO
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: NO
Beginning charge:
Ending charge:
Currently charging:

Opportunity rover used: YES
Hours: 16.8
Beginning charge: 100%
Ending charge: 75%
Currently charging: YES

Curiosity rover used: NO
Beginning charge:
Ending charge:
Currently charging:

HabCar used and why, where : NO

General notes and comments: –

Summary of internet: CapCom Bernard wanted some information about the Hab Router. I gave him it by email.

Summary of suits and radios: Backpack 2 & 6 put aside.

Summary of Hab operations: –

Summary of GreenHab operations: Jeremy (GreenHab officer) began to set the new shade with the help of other crew members.

Summary of ScienceDome operations: –

Summary of RAMM operations: –

Summary of health and safety issues: –

Questions, concerns and requests to Mission Support:

Bulb :
Is it possible to have a new light bulb for the bathroom ?
Brand, model and specifications of the old one : Lights of America – 2814 S – 120V 60Hz 14W 255mA

Pipe leak :
Does a map of all Hab pipes exist ? I still don’t find were the leak comes from. I thought it could be a cooling or heating water, therefore I checked both heaters (roof and wall) today, and they are OK.
The crew is out of ideas considering this leak (I also checked taps yesterday), therefore such a map can help me.

Trashes :
The crew needs indications about trashes to know exactly in which ones burnable and unburnable wastes have to be put into.

Copyright © The Mars Society. All rights reserved. | Main Site